OpenStudy (anonymous):

y^-5y/8 * 16y/2y-10?

OpenStudy (anonymous):

\[\frac{y^{-5}y}{8}\cdot\frac{16y}{2y-10}\] This?

OpenStudy (anonymous):

its suposed to y^2-5y/8 *16/2y-10